Beam Test Meeting Mon. Apr. 30, 2012

From GlueXWiki
Jump to: navigation, search
  1. Update on BCal Test (Andrei & Sascha)


Hall B

- nothing to report


- data taken at 5 pass (5.5 GeV) in all 3 working modes (raw, pulse raw, pulse integral)

- attempted to take time extraction data, and failed, crate required rebooting

- Alex will speak with DAQ experts to see if the problem can be resolved

- no configuration changes on Wed, waiting for time extraction mode to work


- data taken at 5 pass (5.5 GeV) including a voltage scan with a temporary fill for the leaking chiller

- several fills over last few days, see log book for more details

- linked presentation above shows motion of central value of adc spectra (beam energy change of 2.2X) of ~2.5X, suspect difference in temperature between runs is cause

- on Wed will go to 20 degree angle position

- hope for an access in next few days to refill chiller and drop temp to investigate peak motion